# revtex4-1Failed to recognize

### v1

@@ Abstract
We show, for any $n\geq 2$, that there exists quantum correlations obtained from performing $n$ dichotomic quantum measurements in a bipartite Bell scenario, which cannot be reproduced by $n-1$ measurements in any causal theory. That is, it requires any no-signaling theory an unbounded number of measurements to reproduce the predictions of quantum theory. We prove our results by showing that there exists Bell inequalities that have to be obeyed by any no-signaling theory involving only $n-1$ measurements and show explicitly how these can be violated in quantum theory. Finally, we discuss the relation of our work to previous works ruling out alternatives to quantum theory with some kind of bounded degree of freedom and consider the experimental verifiability of our results.
Introduction.—Quantum theory is the most successful and accurately tested description of nature that exists in physics Hanneke et al. (2008); Sailer et al. (2022). In its essence, quantum theory is a set of mathematical rules leading to observable predictions. However, we are still lacking the concise physical principles obeyed by nature that lead to the corresponding mathematical framework Wheeler (1989); Chiribella and Spekkens (2016). Even more crucially, to our current knowledge, there is still room for theories beyond quantum theory that might describe nature more accurately, which are in agreement with well-motivated physical principles, such as causality (by virtue of the no-signaling principle) Popescu and Rohrlich (1992); Janotta (2012). It is therefore critical to study alternative theories that are close to quantum theory and to understand whether they can be ruled out Navascués et al. (2015); Pawłowski et al. (2009); Fritz et al. (2013); Brassard et al. (2006); Renou et al. (2021); Coiteux-Roy et al. (2021); Kleinmann and Cabello (2016); Renou et al. (2021).
@@
The most powerful tool available to rule out alternative descriptions of nature to quantum theory is Bell’s theorem Bell (1964) and Bell’s inequalities. Initially devised to rule out the possibility of theories based on the premise of local hidden variables Einstein et al. (1935), which has been demonstrated in many groundbreaking experiments Rosenfeld et al. (2017); Giustina et al. (2015); Hensen et al. (2015); Shalm et al. (2015), the success of Bell’s theorem goes way beyond its initial use Brunner et al. (2014); Tavakoli et al. (2022). In particular, Bell’s theorem can be used to self-test quantum systems Šupić and Bowles (2020), proof the security of cryptography protocols with minimal assumptions Ekert (1991); Acín et al. (2007), and for witnessing the generation of genuine randomness Colbeck (2009); Pironio et al. (2010); Colbeck and Kent (2011).
On a more fundamental level, the violation of a Bell inequality certifies the presence of entangled systems Horodecki et al. (2009); Moroder et al. (2013) (i.e., systems that cannot be prepared by local operations only) and incompatible measurements Gühne et al. (2023); Heinosaari et al. (2016); Chen et al. (2021); Quintino et al. (2019); Fine (1982); Wolf et al. (2009); Cavalcanti and Skrzypczyk (2016) (i.e., specific observable quantities that cannot be measured jointly via a single measurement). Crucially, entanglement and incompatibility are necessary requirements for the violation of a Bell inequality not only in quantum theory but in much more general theories leading to probabilistic observations, so called general probabilistic theories Acín et al. (2010); Plávala (2023); Banik (2015); Plávala (2016), which quantum theory is a special case of.
@@
Judiciously chosen Bell scenarios have been used in the past to rule of causal theories of nature with only: a finite number of measurement outcomes Kleinmann and Cabello (2016), a finite number of entangled parties Collins et al. (2002); Seevinck and Svetlichny (2002), and sources of correlations of finite degree Coiteux-Roy et al. (2021). Hence, Bell’s theorem can be used to proof that nature does not allow for a finite bound on many external parameters (i.e., parameters that are not part of the mathematical theory but more of the Bell experiment itself).
@@
Figure 1: Graphical representation of our main question. Measurement devices are treated as black-boxes, without even assuming the physical theory they are governed by. Just on the basis of the input-output statics, we rule out that the inner workings of the measurement device include a randomized process $\mathrm{R}$ and measurement devices with fewer inputs to simulate the considered device. While it is here only shown for $3$ measurements that are simulated by devices of effectively $2$ measurements, we rule out that this is generally possibility for all causal theories and any finite number of measurements $n\geq 2$.
@@
However, the question regarding the possible boundlessness of the number of measurements needed to describe nature remains unanswered. Surprisingly, recent progress Kleinmann et al. (2017); Quintino et al. (2019) has been only able to show that nature cannot be described with just $2$ measurements (see also Figure 1), leaving a general answer to the question: Can nature possibly be described by a (maybe very exotic) causal theory that only requires a finite number of measurements? widely open.
@@
In this work, we answer exactly this question in the negative, that means we show that nature cannot be described by any causal theory that involves only a finite number of measurements. In particular, we show the existence of quantum correlations obtained from performing $n$ different dichotomic measurement on an bipartite entangled quantum state that cannot be reproduced by any theory of nature (classical, quantum, and beyond), which is compatible with causality (i.e., it does not allow for super luminal signalling) involving only $n-1$ measurements for every finite $n$. Hence, every possible way to come up with a, maybe very exotic, causal theory of nature in which there is an (arbitrarily large but finite) upper limit on the number of different measurement settings, can be falsified. We want to emphasize at this point that the terms causal theory or causality, which are overloaded with various non-equivalent definitions in quantum information theory, are used synonymous to no-signaling theory throughout this manuscript.
@@
To prove our result, we consider the theory of measurement incompatibility, whose extension to recently studied incompatibility structures Kunjwal et al. (2014); Heinosaari et al. (2008); Liang et al. (2011); Quintino et al. (2019); Tendick et al. (2023) is crucial to formally define what it means for $n$ measurements to be genuinely incompatible, i.e., not reducible to $n-1$ measurements. We then connect genuine incompatibility to the family of $M_{nn22}$ inequalities Brunner et al. (2006), to show that a violation of these inequalities implies the existence of genuinely $n$ measurements in a theory independent way, i.e., without assuming the formalism of a specific theory but only relying on causality by virtue of the no-signaling principle. We then show that any of the $M_{nn22}$ can be violated by quantum theory with a quantum state of local dimension $n$. Finally we discuss possible experimental implementations of our findings.
@@
Incompatible measurements in nature.—Let us start by reviewing the notions of incompatibility and how to test device and theory independently that there exist observable quantities that cannot be measured jointly. In quantum theory, measurements are described by positive semidefinite operators $A_{a|x}\geq 0$ acting on a Hilbert space $\mathcal{H}\cong\mathds{C}^{d}$ of dimension $d$, such that the completeness relation $\sum_{a}A_{a|x}=\mathds{1}$, (where $\mathds{1}$ is the identity operator), is fulfilled. Here, $x$ denotes the label of a specific measurement and $a$ the label of the corresponding outcome to that measurement. We say a set of measurements (also known as assemblage) $\mathcal{A}=\{\{A_{a|x}\}_{a}\}_{x}$ is jointly measurable if it can be simulated by performing a single measurement $\{G_{\lambda}\}_{\lambda}$ (the so-called parent positive operator valued measure (POVM)) and some classical post-processing via the conditional probabilities $p(a|x,\lambda)$ such that
$A_{a|x}=\sum_{\lambda}p(a|x\lambda)G_{\lambda},$ (1)
@@
holds. An assemblage that is not jointly measurable is called incompatible. Note that joint measurability can conceptually be seen as the generalization of commutativity of projective measurements to general quantum measurements Heinosaari et al. (2016); Gühne et al. (2023).
@@
To demonstrate incompatibility of measurements without needing a full characterization of the measurement device, or the physical theory it is governed by, one needs to rely on the device-independent paradigm and the violation of Bell inequalities. Let us consider a bipartite Bell experiment in which two distant parties, Alice and Bob, perform one out of $n$ measurements in each round of the experiment labeled by $x$ and $y$, respectively. In each round, Alice obtains one out of two outcomes $a$, similarly Bob obtains outcome $b$. Treating their devices as black-boxes, all information about the experiment that is accessible to Alice and Bob are the input-output statistics, also called behaviour or correlations, $\mathbf{P}=\{p(ab|xy)\}$. Just on the basis of the obtained behaviour $\mathbf{P}$ we want to prove that the considered black-box devices involve indeed incompatible measurements and hence cannot be treated as a single measurement.
@@
By definition of being a Bell inequality, the CHSH inequality has to be obeyed by all local (also known as classical) theories, i.e., by all theories where the correlations $\mathbf{P}$ could be explained via pre-shared randomness. However, it is well-known that quantum correlations obtained from performing quantum measurements on a shared bipartite quantum state $\rho$ such that $p(ab|xy)=\mathrm{Tr}[(A_{a|x}\otimes B_{b|y})\rho]$ can violate the CHSH inequality, provided judiciously chosen incompatible measurements $\mathcal{A},\mathcal{B}$ are performed on an entangled state $\rho$. That is, the CHSH inequality certifies in a device-independent way the existence of entanglement and for us more importantly, that Alice and Bob each employed incompatible measurements. While incompatibility is necessary for the violation of a Bell inequality, it is known that is generally not sufficient Bene and Vértesi (2018); Hirsch et al. (2018) and it is an open question which incompatibility structures can be certified device-independently.
@@
In their seminal work, Wolf et al. Wolf et al. (2009) prove that any pair of incompatible dichotomic measurements can be used to violate the CHSH inequality if the measurements of the other party and the quantum state $\rho$ are optimized accordingly. Moreover, the authors in Wolf et al. (2009) prove that the violation of a general Bell inequality not only implies the incompatibility of measurements in quantum theory, but in any theory that obeys the no-signaling principle. The no-signaling principle states that the measurement devices of Alice and Bob cannot be used for superluminal signaling, which can be seen as a basic definition for a physical theory to obey causality. More formally, the correlations $\mathbf{P}$ obey the no-signaling principle Popescu and Rohrlich (1992); Janotta (2012), if their marginal distributions $p_{A}(a|x)$ and $p_{B}(b|y)$ are well-defined and independent of the input of the other party, i.e., $p_{A}(a|x)=p_{A}(a|xy)=\sum_{b}p(ab\lvert xy)$ for all $a,x,$ and similarly for Bob’s marginal. However, the general question of whether there always exist $n$ measurements, for any $n\geq 2$, that cannot be reproduced by at most $n-1$ measurements in any causal theory remains still open.
@@
Boundlessly many genuine incompatible measurements in nature.—Here, we answer this question affirmatively. To hat end, we first have to define properly what it means to have genuinely $n$ measurements. While ordinary incompatibility according to Eq. (1) only guarantees that the assemblage $\mathcal{A}$ cannot be reduced to a single measurement, it does not guarantee that $n$ measurements cannot be reduced to fewer measurements in general. Using the notion of recently studied incompatibility structures, the authors in Quintino et al. (2019) define an assemblage to contain (at least) genuinely $n$ measurements if it cannot be implemented by probabilistically employing assemblages that contain at most genuinely $n-1$ measurements. That is, if the assemblage $\mathcal{A}$ cannot be decomposed such that
@@
where $\mathcal{F}^{(s,t)}\in\mathrm{JM}^{(s,t)}$ are assemblages, employed with probability $p_{(s,t)}$, in which the pair $(s,t\neq s)$ is jointly measurable, which makes each $\mathcal{F}^{(s,t)}$ contain at most genuinely $n-1$ measurements (as the pair $(s,t\neq s)$ is one effective measurement) and the set $\mathcal{T}$ contains all tuples $(s,t\neq s)$ constructed from the integers $\{1,\cdots,n\}$.
@@
where $p^{(s,t)}(ab|xy)$ describes a behaviour $\mathbf{P}^{(s,t)}$ that is local once it is restricted to only the measurements $(s,t)$ on Bob’s side. Now, crucially, disproving that a decomposition according to Eq. (4) exists while only assuming the no-signaling principle, constitutes a device and theory independent proof of having access to genuinely $n$ measurements.
The numerical analysis for the case $n=3$ in Quintino et al. (2019) reveals that while it is easy to use many well known inequalities like the $3$-CHSH game Bowles et al. (2018), the elegant Gisin (2007), or the chained Bell inequality Braunstein and Caves (1990) for the certification of genuinely $3$ measurements assuming quantum theory, it is more difficult to rule out the possibility of a simulation by only $2$ measurements in general no-signaling theories. Essentially, there exists only one relevant Bell inequality, the $M_{3322}$ inequality for this task (see also Kleinmann et al. (2017)), which was already and independently been considered in Brunner et al. (2004). However, the $M_{3322}$ inequality only allows for a relatively small violation within quantum theory and employing numerical methods can only provide insights for small number of measurements $n$. Hence, despite providing the best current progress, the methods in Bowles et al. (2018) are of limited use for answering our main question.
We show in the following that a detailed analytical analysis of why the $M_{3322}$ certifies genuinely $3$ measurements in any causal theory leads to the key observation that enables us to prove our main result. More precisely, it enables us for every finite integer $n\geq 2$, to propose a Bell experiment, including $n$ dichotomic quantum measurements, that cannot be reproduced by any causal (i.e., no-signaling) theory using only $n-1$ measurements. Let us denote the set of no-signaling correlations of at most genuinely $n$ measurements according to Eq. (4) by $\mathrm{GNS}_{n}$.
@@
For every $n\geq 2$ there exist quantum correlations $\mathbf{Q}\in\mathrm{GNS}_{n}$ that cannot be reproduced by any no-signaling correlations with at most genuinely $n-1$ measurements, i.e., $\mathbf{Q}\notin\mathrm{GNS}_{n-1}$.
@@
Let us start by analytically analyzing why the $M_{3322}$ inequality cannot be violated by distributions $\mathbf{P}\in\mathrm{GNS}_{2}$. To make the proof as vivid as possible, we will work with the Collins-Gisin table notation of behaviours $\mathbf{P}$ and Bell inequalities to understand the structure of the considered inequalities better. In particular, any normalized no-signaling behaviour $\mathbf{P}$ of two-outcome measurements in uniquely characterized by a table (shown here for two settings)
$\begin{array}[]{c||c c}&P(A_{1})&P(A_{2})\\ \hline\cr\hline\cr P(B_{1})&P(A_{1}B_{1})&P(A_{2}B_{1})\\ P(B_{2})&P(A_{1}B_{2})&P(A_{2}B_{2})\\ \end{array},$ (8)
@@
Using the fact that for any no-signaling distribution the marginal distributions upper bound their involved correlations, i.e., $P(A_{k})\geq P(A_{k}B_{j})$ and $P(B_{k})\geq P(A_{j}B_{k})$, we can deduce consequences on the incompatibility structures that follow from a violation of $M_{3322}$. Namely, it follows that $M_{3322}>0$ implies a simultaneous violation of the inequalities
@@
To see this, we only have to realize that simultaneously erasing positive contributions like $P(A_{k}B_{j})$ and negative contributions $P(A_{k})$ makes it easier to violate the inequality, as the total contribution of these terms would be negative. A similar reasoning follows for erasing terms $P(A_{k}B_{j})$ that contribute with a negative sign. Now, this means every pair of two measurements of Bob has to lead to nonlocal correlations, as each of the above versions of the $M_{3322}^{(s,t)}$ is equivalent to the CHSH inequality for the pair $(s,t)$. Hence, every pair of Bob’s measurements has to be incompatible. By contra position, if one the pairs $(s,t)$ is compatible, it holds $M_{3322}^{(s,t)}\leq 0$. That means, the upper bound $M_{3322}\leq 0$ holds for all no-signaling distributions $\mathbf{P}$ in which some pair $(s,t)$ leads to local correlations. Now, by convexity of the decomposition in Eq. (4) and the fact that $M_{3322}$ is a linear Bell inequality, the same also holds for any behaviour $\mathbf{P}\in\mathrm{GNS}_{2}$.
For general $n\geq 2$, we consider the $M_{nn22}$ inequalities Brunner et al. (2006), (of which $M_{3322}$ is simple the $n=3$ case) derived from slightly varying the family of $I_{nn22}$ inequalities Collins and Gisin (2004). More formally, it holds $M_{nn22}\coloneqq I_{nn22}-(n-2)p_{A}(a=0|x=1)$ or
@@
We generalize the above analysis in the Supplemental Material (SM) Sup , i.e., we show that there exist a version of the CHSH inequality for every pair $(s,t)$ of Bob’s measurements that has to be violated if $M_{nn22}$ is violated. Consequently, it holds $M_{nn22}\leq 0$ for all correlations $\mathbf{P}\in\mathrm{GNS}_{n-1}$ originating from at most $n-1$ genuine measurements on Bob’s side according to Eq. (4).

### v2

@@ Abstract
We show, for any finite $n\geq 2$, that there exist quantum correlations obtained from performing $n$ dichotomic quantum measurements in a bipartite Bell scenario, which cannot be reproduced by mixtures of measurement devices with at most $(n-1)$ incompatible measurements across different partitions in any no-signaling theory. That is, it requires any no-signaling theory an unbounded number of measurements to reproduce the predictions of quantum theory. We prove our results by showing that there exist linear Bell inequalities that have to be obeyed by any no-signaling theory involving only $(n-1)$-wise incompatible measurements and show explicitly how these can be violated in quantum theory. Finally, we discuss the relation of our work to previous works ruling out alternatives to quantum theory with some kind of bounded degree of freedom and consider the experimental verifiability of our results.
Introduction.—The goal of theoretical physics is to devise a model of nature that accurately explains past observations and allows for experimentally verifiable predictions. Quantum theory is the most successful and accurately tested description of nature that exists in physics Hanneke et al. (2008); Sailer et al. (2022). Despite $100$ years of proving its success, the effectiveness of quantum theory is still puzzling from a theory-building perspective Wheeler (1989); Chiribella and Spekkens (2016). In its essence, quantum theory is a set of mathematical rules that lead to observable predictions.
Notably, several works indeed derive the mathematical formalism from first principles Hardy (2001); Chiribella et al. (2011); Barnum et al. (2014); Cabello (2019); Dakic and Brukner (2009); Masanes and Müller (2011). However, the search for the precise physical principles Popescu and Rohrlich (1992); Pawłowski et al. (2009); Fritz et al. (2013); Brassard et al. (2006); Linden et al. (2007); Navascués and Wunderlich (2009); Navascués et al. (2015) that allow for a derivation of (the limits of) quantum theory purely from observations, i.e., the input-output statistics of local measurements in a Bell scenario with uncharacterized devices (black-boxes) is still ongoing.
@@
Initially devised to rule out the possibility of theories based on the premise of local hidden variables Einstein et al. (1935), which has been demonstrated in many groundbreaking experiments Rosenfeld et al. (2017); Giustina et al. (2015); Hensen et al. (2015); Shalm et al. (2015), Bell’s theorem Bell (1964) and Bell inequalities found many applications beyond their initial use Brunner et al. (2014); Tavakoli et al. (2022). The violation of a Bell inequality certifies the presence of entangled systems Horodecki et al. (2009); Moroder et al. (2013) and incompatible measurements Gühne et al. (2023); Heinosaari et al. (2016); Chen et al. (2021); Quintino et al. (2019); Fine (1982); Wolf et al. (2009); Cavalcanti and Skrzypczyk (2016). Moreover, Bell inequality violations are a necessary prerequisite to self-test quantum systems Šupić and Bowles (2020), to prove the security of cryptography protocols with minimal assumptions Ekert (1991); Acín et al. (2007), and to witness the generation of randomness Colbeck (2009); Pironio et al. (2010); Colbeck and Kent (2011).
Crucially, entanglement and incompatibility are necessary requirements for the violation of a Bell inequality not only in quantum theory but also in much more general theories leading to probabilistic observations, so-called general probabilistic theories Acín et al. (2010); Plávala (2023); Banik (2015); Plávala (2016), which quantum theory is a special case of. Moreover, alternative theories of nature that are close to quantum theory Renou et al. (2021); Coiteux-Roy et al. (2021); Kleinmann and Cabello (2016); Collins et al. (2002); Seevinck and Svetlichny (2002) can be ruled-out through the quantum violation of Bell-like inequalities restricting correlations obeyed by these theories.
@@
Judiciously chosen Bell scenarios have been used in the past to rule out no-signaling theories reproducing quantum correlations with only: a finite number of measurement outcomes Kleinmann and Cabello (2016), a finite number of entangled parties Collins et al. (2002); Seevinck and Svetlichny (2002), and sources of correlations of finite degree Coiteux-Roy et al. (2021). Hence, Bell-like inequalities can be used to prove that quantum correlations do not allow for a finite bound on many external parameters (i.e., parameters that are not part of the mathematical theory but more of the Bell experiment itself).
@@
Figure 1: Graphical representation of our main question for the special case of $3$ measurements. Measurement devices are treated as black-boxes, i.e., we only observe the input-output statistics of the devices. Moreover, we only assume the devices to be restricted by the no-signaling principle and do not refer specifically to the quantum formalism. Just on the basis of the input-output statistics, we rule out that the inner workings of the measurement device include a randomization process $\mathrm{R}$ and measurement devices with fewer inputs to simulate the considered device.
@@
However, the question regarding the possible boundlessness of the number of measurements needed to describe quantum correlations remains unanswered. Surprisingly, recent progress Kleinmann et al. (2017); Quintino et al. (2019) has been only able to show that quantum theory cannot be reproduced with just $2$ measurements (see also Figure 1), leaving a general answer to the question: Can quantum theory possibly be reproduced by a (maybe very exotic) no-signaling theory that only requires a finite number of measurements? widely open.
@@
In this work, we answer exactly this question in the negative. In particular, we show the existence of quantum correlations obtained from performing $n$ different dichotomic measurements on an bipartite entangled quantum state that cannot be reproduced by any theory (classical, quantum, and beyond), which is compatible with the no-signaling principle involving only $(n-1)$ measurements for every finite $n$. Hence, every possible way to come up with a no-signaling theory reproducing quantum theory in which there is an (arbitrarily large but finite) upper limit on the number of different measurement settings, can be falsified.
@@
To prove our result, we consider the theory of measurement incompatibility, whose extension to recently studied incompatibility structures Kunjwal et al. (2014); Heinosaari et al. (2008); Liang et al. (2011); Quintino et al. (2019); Tendick et al. (2023) is crucial to formally define what it means to perform genuinely $n$ different measurements. More precisely, we use that measurements which are genuinely $n$-wise incompatible, cannot be reduced to effectively $(n-1)$ measurements as convex combinations of assemblages that are at most $(n-1)$-wise incompatible. We then connect genuine $n$-wise incompatibility to the family of $M_{nn22}$ inequalities Brunner et al. (2006), to show that a violation of these inequalities implies the existence of genuinely $n$-wise incompatible measurements in any no-signaling theory. We then show that any of the $M_{nn22}$ can be violated by quantum theory with a quantum state of local dimension $n$. Finally, we discuss possible experimental implementations of our findings.
@@
Incompatible measurements in Bell experiments.—Let us start by reviewing the notions of measurement incompatibility. In quantum theory, measurements are described by positive semidefinite operators $A_{a|x}\geq 0$ acting on a Hilbert space $\mathcal{H}\cong\mathds{C}^{d}$ of dimension $d$, such that the completeness relation $\sum_{a}A_{a|x}=\mathds{1}$, (where $\mathds{1}$ is the identity operator), is fulfilled. Here, $x$ denotes the label of a specific measurement and $a$ the label of the corresponding outcome to that measurement. We say a set of measurements (also known as assemblage) $\mathcal{A}=\{\{A_{a|x}\}_{a}\}_{x}$ is jointly measurable if it can be simulated by performing a single measurement $\{G_{\lambda}\}_{\lambda}$ (the so-called parent positive operator valued measure (POVM)) and some classical post-processing via the conditional probabilities $p(a|x,\lambda)$ such that
$A_{a|x}=\sum_{\lambda}p(a|x,\lambda)G_{\lambda},$ (1)
@@
holds. An assemblage that is not jointly measurable is called incompatible. Conceptually, joint measurability can be seen as the generalization of commutativity of projective measurements to general quantum measurements Heinosaari et al. (2016); Gühne et al. (2023).
@@
To certify the incompatibility of measurements without characterizing the measurement device, one can rely on the device-independent (black-box) paradigm and the violation of Bell inequalities. Let us consider a bipartite Bell experiment in which two distant parties, Alice and Bob, perform one out of $n$ measurements in each round of the experiment labeled by $x$ and $y$, respectively. In each round, Alice obtains one out of two outcomes $a$, similarly Bob obtains outcome $b$. Treating their devices as black-boxes, all information about the experiment that is accessible to Alice and Bob are the input-output statistics, also called behavior or correlations, $\mathbf{P}=\{p(ab|xy)\}$. Just on the basis of the obtained behaviour $\mathbf{P}$ we want to prove that the considered black-box devices involve indeed incompatible measurements and hence cannot be treated as a single measurement.
@@
By definition of being a Bell inequality, the CHSH inequality has to be obeyed by all local (also known as classical) theories, i.e., by all theories where the correlations $\mathbf{P}$ could be explained via pre-shared randomness. However, it is well-known that quantum correlations obtained from performing quantum measurements on a shared bipartite quantum state $\rho$ such that $p(ab|xy)=\mathrm{Tr}[(A_{a|x}\otimes B_{b|y})\rho]$ can violate the CHSH inequality, provided judiciously chosen incompatible measurements $\mathcal{A},\mathcal{B}$ are performed on an entangled state $\rho$. That is, the CHSH inequality constitutes a device-independent certificate of entanglement and measurement incompatibility. While incompatibility is necessary for the violation of a Bell inequality, it is known that is generally not sufficient Bene and Vértesi (2018); Hirsch et al. (2018), however all incompatibility structures can be certified device-independently Yadavalli et al. (2024).
@@
In their seminal work, Wolf et al. Wolf et al. (2009) prove that the violation of any Bell inequality (and in particular the CHSH inequality) not only implies the incompatibility of measurements in quantum theory, but in any theory that obeys the no-signaling principle. The no-signaling principle states that the measurement devices of Alice and Bob cannot be used for superluminal signaling, which means Alice is oblivious of Bob’s measurement setting and vice versa. More formally, the correlations $\mathbf{P}$ obey the no-signaling principle Popescu and Rohrlich (1992); Janotta (2012), if their marginal distributions $p_{A}(a|x)$ and $p_{B}(b|y)$ are well-defined and independent of the input of the other party, i.e., $p_{A}(a|x)=p_{A}(a|xy)=\sum_{b}p(ab\lvert xy)$ for all $a,x,$ and similarly for Bob’s marginal. However, the general question of whether there always exist $n$ quantum measurements, for any $n\geq 2$, that cannot be reproduced by at most $(n-1)$ measurements in any no-signaling theory remains still open.
@@
Boundlessly many genuine incompatible measurements.—Here, we answer this question affirmatively. To that end, we first have to define properly what it means to perform genuinely $n$ different measurements through the notion of genuinely $n$-wise incompatible measurements. While ordinary incompatibility according to Eq. (1) only guarantees that the assemblage $\mathcal{A}$ cannot be reduced to a single measurement, it does not guarantee that $n$ measurements cannot be reduced to fewer measurements in general. Using the notion of recently studied incompatibility structures, the authors in Quintino et al. (2019) define an assemblage to contain (at least) genuinely $n$-wise incompatible measurements if it cannot be implemented by probabilistically employing assemblages that contain at most genuinely $(n-1)$-wise incompatible measurements. That is, if the assemblage $\mathcal{A}$ cannot be decomposed such that
@@
where $\mathcal{F}^{(s,t)}\in\mathrm{JM}^{(s,t)}$ are assemblages, employed with probability $p_{(s,t)}$, in which the pair $(s,t\neq s)$ is jointly measurable. This makes each $\mathcal{F}^{(s,t)}$ contain at most genuinely $(n-1)$-wise incompatible measurements (as the pair $(s,t\neq s)$ is one effective measurement) and the set $\mathcal{T}$ contains all tuples $(s,t\neq s)$ constructed from the integers $\{1,\cdots,n\}$.
@@
where $p^{(s,t)}(ab|xy)$ describes a behaviour $\mathbf{P}^{(s,t)}$ that is local once it is restricted to only the measurements $(s,t)$ on Bob’s side. Now, crucially, disproving that a decomposition according to Eq. (4) exists while only assuming the no-signaling principle, constitutes a device-independent proof of genuinely $n$-wise incompatible measurements in any no-signaling theory.
The numerical analysis for the case $n=3$ in Quintino et al. (2019) reveals that while it is easy to use many well known inequalities like the $3$-CHSH game Bowles et al. (2018), the elegant Gisin (2007), or the chained Bell inequality Braunstein and Caves (1990) for the certification of genuinely $3$-wise measurements assuming quantum theory, it is more difficult to rule out the possibility of a simulation by only $2$ measurements in general no-signaling theories. Essentially, there exists only one relevant Bell inequality, the $M_{3322}$ inequality for this task (see also Kleinmann et al. (2017)), which has also already been considered in Brunner et al. (2004). However, the $M_{3322}$ inequality only allows for a relatively small violation within quantum theory and employing numerical methods can only provide insights for small number of measurements $n$. Hence, despite providing the best current progress, the methods in Bowles et al. (2018) are of limited use for answering our main question.
We show in the following that a detailed analytical analysis of why the $M_{3322}$ certifies genuinely $3$-wise incompatible measurements in any no-signaling theory leads to the key observation that enables us to prove our main result. More precisely, it enables us for every finite integer $n\geq 2$, to propose a Bell experiment, including $n$ dichotomic quantum measurements, that cannot be reproduced by any no-signaling theory using only $(n-1)$-wise compatible measurements. Let us denote the set of no-signaling correlations of at most genuinely $(n-1)$-wise incompatible measurements by $\mathrm{GNS}_{n-1}$. That is, $\mathbf{P}\in\mathrm{GNS}_{n-1}$ if a decomposition according to Eq. (4) exists.
@@
For every $n\geq 2$ there exist quantum correlations $\mathbf{Q}$ obtained from performing $n$ quantum measurements that cannot be reproduced by any no-signaling correlations with at most genuinely $(n-1)$-wise incompatible measurements, i.e., $\mathbf{Q}\notin\mathrm{GNS}_{n-1}$.
@@
Let us start by analytically analyzing why the $M_{3322}$ inequality cannot be violated by distributions $\mathbf{P}\in\mathrm{GNS}_{2}$. To make the proof as vivid as possible, we will work with the Collins-Gisin table notation of behaviours $\mathbf{P}$ and Bell inequalities to understand the structure of the considered inequalities better. In particular, any normalized no-signaling behaviour $\mathbf{P}$ of two-outcome measurements in uniquely characterized by a table (shown here for three settings)
$\begin{array}[]{c||c c c}&P(A_{1})&P(A_{2})&P(A_{3})\\ \hline\cr\hline\cr P(B_{1})&P(A_{1}B_{1})&P(A_{2}B_{1})&P(A_{3}B_{1})\\ P(B_{2})&P(A_{1}B_{2})&P(A_{2}B_{2})&P(A_{3}B_{2})\\ P(B_{3})&P(A_{1}B_{3})&P(A_{2}B_{3})&P(A_{3}B_{3})\end{array},$ (9)
@@
Now, as we are restricting ourselves only to the no-signaling principle, consider a no-signaling distribution $\mathbf{P}$ such that $M_{3322}\odot\mathbf{P}>0$, where the symbol $\odot$ denotes the application of the Bell functional onto the distribution $\mathbf{P}$. Note that for any no-signaling distribution the marginal distributions upper bound their involved correlations, i.e., $P(A_{k})\geq P(A_{k}B_{j})$ and $P(B_{j})\geq P(A_{k}B_{j})$. From this, we can deduce that
$[P(A_{2}B_{1})-P(B_{1})]+[P(A_{2}B_{2})-P(B_{2})]$
$+[P(A_{1}B_{3})-P(A_{1})]-P(A_{2}B_{3})\leq 0,$ (15)
for any-no signaling distribution $\mathbf{P}$. Note that we also used here that probabilities, such as the term $P(A_{2}B_{3})$ are non-negative. Hence, it follows that $M_{3322}\odot\mathbf{P}>0$ implies also $M_{3322}^{(1,2)}\odot\mathbf{P}>0$, where $M_{3322}^{(1,2)}$ is given by
@@
as the term in Eq. (15) only contributes non-positively towards the $M_{3322}$ inequality. Note that $M_{3322}^{(1,2)}$ is just the CHSH inequality involving Bob’s measurements $(s=1,t=2)$. By noting that other terms that only contribute negatively towards the Bell score of the $M_{3322}$ can be eliminated similarly, we also deduce that $M_{3322}\odot\mathbf{P}>0$ implies a violation of the inequalities
@@
Now, this means every pair of two measurements of Bob has to lead to nonlocal correlations, as each of the above versions of the $M_{3322}^{(s,t)}$ is equivalent to the CHSH inequality for the pair $(s,t)$. Hence, every pair of Bob’s measurements has to be incompatible. By contra position, if one the pairs $(s,t)$ is compatible, it holds $M_{3322}^{(s,t)}\leq 0$. Consequently, it holds $M_{3322}\leq 0$ for all no-signaling distributions $\mathbf{P}$ in which some pair $(s,t)$ is compatible. The final question is, whether a behavior $\mathbf{P}$ s.t. $M_{3322}\odot\mathbf{P}>0$ could still be written as a convex combination of different $\mathbf{P}^{(s,t)}$ according to Eq. (4). However, using the fact that $M_{3322}$ is a linear in the probabilities and Eq. (4) is a convex decomposition, it follows directly that that $M_{3322}\odot\mathbf{P}\leq 0$ for any behavior $\mathbf{P}\in\mathrm{GNS}_{2}$.
For general $n\geq 2$, we consider the $M_{nn22}$ inequalities Brunner et al. (2006), derived from slightly varying the family of $I_{nn22}$ inequalities Collins and Gisin (2004). More formally, it holds $M_{nn22}\coloneqq I_{nn22}-(n-2)p_{A}(a=0|x=1)$ or
@@
We show the full proof for general $n$, generalizing the above analysis in the Appendix. More specifically, we show that there exist a version of the CHSH inequality for every pair $(s,t)$ of Bob’s measurements that has to be violated if $M_{nn22}$ is violated. This means that a distribution $\mathbf{P}$ that violates $M_{nn22}$ will exhibit nonlocality (and hence incompatibility) in each pair $(s,s\neq t)$. Again, by the convexity of the problem and using the fact that $M_{nn22}$ is linear, it holds $M_{nn22}\odot\mathbf{P}\leq 0$ for all correlations $\mathbf{P}\in\mathrm{GNS}_{n-1}$ originating from at most $(n-1)$-wise genuinely incompatible measurements on Bob’s side according to Eq. (4).